OpenStudy (august899):

Which expressions are equivalent to 1+sinθ/1-sinθ ? Check all that apply. (Multiple Choice)

OpenStudy (august899):

A. cos^2θ/(1+sinθ)^2 B. cos^2θ/(1-sinθ)^2 C.( 1+sinθ)^2/cosθ^2 D. (1-sinθ)^2/cosθ^2

OpenStudy (anonymous):

multiply by \[\frac{ 1-\sin \theta }{ 1-sin \theta }\]

OpenStudy (august899):

Ok. Is it D?